OpenStudy (anonymous):

Use substitution to solve the system. x=2y+8 2x-3y=13

OpenStudy (dmezzullo):

erm, multiply the first equation by \(-2\). that yields\[\begin{cases}-2x = -4y - 16 \\ 2x - 3y = 13\end{cases}\]add both equations and solve for \(y\).

OpenStudy (johnweldon1993):

you already know what x equals...so sub it in for x in the second equation 2(2y+8) - 3y = 13 Now just distribute and simplify

OpenStudy (johnweldon1993):

dmezz thats elimination ...the problem asked for substitution :)

OpenStudy (anonymous):

solve each system by the elimination method 1)3x+2y=10 2x+y=7 2)3x-2y=7 2x-y=8 3)2x+3y=-13 -4x+6y=-10

OpenStudy (harsimran_hs4):

you already have the value of x from the first equation just substitute that value and get the answer

OpenStudy (dmezzullo):

ah, yes. my bad